OpenStudy (anonymous):

One skier covered the distance of 20km in 20 min faster than the other skier. Find the speed of both skiers if one of them is travelling at 2km/hour faster than the other one.

OpenStudy (anonymous):

@jim_thompson5910 @Mertsj PLEASE HELP ME

jimthompson5910 (jim_thompson5910):

let x = speed of slower skier y = speed of faster skier

jimthompson5910 (jim_thompson5910):

"one of them is travelling at 2km/hour faster than the other one.", so by this statement and our variable definitions, we can say y = x+2

jimthompson5910 (jim_thompson5910):

now use d = rt to set up a distance/rate/time equation for the slower skier d = rt 20 = x*t solve for t to get t = 20/x and do the same for the faster skier d = rt 20 = y*(t-20) 20 = y*(20/x-20) ... plug in t = 20/x 20 = (x+2)*(20/x-20) ... plug in y = x+2 now solve for x